$A$ bag contains $3$ red balls and $5$ black balls. $A$ ball is drawn at random from the bag. What is the probability that the ball drawn is
$(i)$ red?
$(ii)$ not red?

Vedclass pdf generator app on play store
Vedclass iOS app on app store
(A) $(i)$ Total number of balls in the bag $= 3 + 5 = 8$.
The probability of drawing a red ball is given by the formula: $P(\text{red}) = \frac{\text{Number of favourable outcomes}}{\text{Total number of possible outcomes}}$.
$P(\text{red}) = \frac{3}{8}$.
$(ii)$ The probability of not drawing a red ball is the complement of drawing a red ball.
$P(\text{not red}) = 1 - P(\text{red})$.
$P(\text{not red}) = 1 - \frac{3}{8} = \frac{8 - 3}{8} = \frac{5}{8}$.

Explore More

Similar Questions

$A$ lot consists of $48$ mobile phones of which $42$ are good,$3$ have only minor defects,and $3$ have major defects. Varnika will buy a phone if it is good,but the trader will only buy a mobile if it has no major defect. One phone is selected at random from the lot. What is the probability that it is
$(i)$ acceptable to Varnika?
$(ii)$ acceptable to the trader?

Which of the following arguments are correct and which are not correct? Give reasons for your answer.
$(i)$ If two coins are tossed simultaneously,there are three possible outcomes - two heads,two tails,or one of each. Therefore,for each of these outcomes,the probability is $\frac{1}{3}$.
$(ii)$ If a die is thrown,there are two possible outcomes - an odd number or an even number. Therefore,the probability of getting an odd number is $\frac{1}{2}$.

$(i)$ $A$ lot of $20$ bulbs contains $4$ defective ones. One bulb is drawn at random from the lot. What is the probability that this bulb is defective?
$(ii)$ Suppose the bulb drawn in $(i)$ is not defective and is not replaced. Now one bulb is drawn at random from the rest. What is the probability that this bulb is not defective?

Two balanced dice are rolled. Find the probability that the sum of numbers on two dice lies between $5$ and $9$.

Cards with numbers $2$ to $101$ are placed in a box. $A$ card is selected at random. Find the probability that the card has
$(i)$ an even number
$(ii)$ a square number

Vedclass Products

For Students

Vedclass Test Series

Mock tests in real JEE/NEET style with performance analysis. 5-day free trial.

Start Free Trial
For Teachers

Exam Paper Generator

Generate Set A/B/C/D exam papers from 7.5L+ questions in 2 minutes. 3 chapters free.

Try Free
For Institutes

Online Exam Module

Live online exams with unlimited students, 360° analytics & white-label branding.

See Demo